OpenStudy (anonymous):

evaluate the expression x+(3x)^2 +x where x =the square of 2 divided by 2

OpenStudy (anonymous):

where x = the square of 2

OpenStudy (anonymous):

x+(3x) 2 +x over 2

OpenStudy (anonymous):

what does X equal...you just said two different things D:

OpenStudy (anonymous):

x = the square of 2\[\sqrt{2}\]

OpenStudy (anonymous):

the sqrt ok \[\frac{ x + 3x^2 + x }{ 2 }\] that it?

OpenStudy (anonymous):

yes!! i couldn't find how to make it do that

OpenStudy (anonymous):

sigh,, only there are parenthesis around the 3x^2

OpenStudy (anonymous):

equation button, anyway put in sqrt of 2 everywhere u see x so \[\frac{ \sqrt{2} + 3\sqrt{2}^2 +\sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

\[(3\sqrt{2}^2) or (3\sqrt{2})^2\]

OpenStudy (anonymous):

yes i got that far

OpenStudy (anonymous):

*\[x+ (3 \sqrt{2})^{2} +x\]

OpenStudy (anonymous):

oh ok that's easy \[\frac{ \sqrt{2} + (3\sqrt{2})^2 +\sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

not easy to me :'(

OpenStudy (anonymous):

but that is the correct way to write it

OpenStudy (anonymous):

well what's sqrt of 2 plus sqrt of 2

OpenStudy (anonymous):

the square root of 4

OpenStudy (anonymous):

which is! lol

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

so i have 8/2 which = 4,,, so my answer is 9 + the square root of 2!!!

OpenStudy (anonymous):

well with that sqrt of 2 it's squared so the square root and the square cancel each other leaving us 2

OpenStudy (anonymous):

got it!!!! thank you!!!!

OpenStudy (anonymous):

\[\frac{ 2 + 9 + 2 }{ 2 }\]

OpenStudy (anonymous):

tyty!!! this makes sense now!!!

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

depends on with what

OpenStudy (anonymous):

5/n-12=-1/n

OpenStudy (anonymous):

don't post it here

OpenStudy (anonymous):

Its me from before I still need help. where?

OpenStudy (anonymous):

Its me from before I still need help. where?

OpenStudy (anonymous):

Where do I post it ?

OpenStudy (anonymous):

I bumped it.

OpenStudy (anonymous):

Please I really need help